• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Olivier

    Log entry time 09:51:18 on June16,2000

    Entry number 43919

    keyword=HP-heater alarm limits changed

    The beam has been off for 45 minutes now (spot change), and the screwed-up HP-heater readout of the target is driving me crazy, because it keeps oscillating between values outside of the limits. I called Joe (cryo on call), who allows me to change this limits to very high values (10000), so that the alarm doesn't trip any more. The value we read on this censor is meaningless anyway !